# Projective geodesic extensions by conformal modifications in nonholonomic mechanics

### v1

@@ 6.4 A mathematical example
Based on the classification in Proposition 12 we know that $(A^{\prime})^{G}$ is a weaker condition than the $\varphi$-simplicity property, but a stronger condition than having only an invariant volume form. To show that the item $(ii)$ of the classification is not void, we give a (mathematical) example of a nonholonomic system that is not $\varphi$-simple, but that does have a solution $f$ of condition $(A^{\prime})^{G}$. Such a system will allow for a projective geodesic extension (with $\overline{g}_{ai}=-G_{ai}$) and an invariant volume form, without being $\varphi$-simple.
@@
$=$ $\frac{1}{2}({\tilde{v}_{x}}^{2}+{\tilde{v}_{y}}^{2}+{\tilde{v}_{z}}^{2}+2{\tilde{v}_{x}}{\tilde{v}_{y}}+2{\tilde{v}_{x}}{\tilde{v}_{z}}+2{\tilde{v}_{y}}{\tilde{v}_{z}}+2{\tilde{v}_{x}}{\tilde{v}_{u}}+2{\tilde{v}_{y}}{\tilde{v}_{u}}+2{\tilde{v}_{z}}{\tilde{v}_{u}}+4{\tilde{v}_{u}}^{2}),$
@@
$G_{bu}R_{ac}^{u}-G_{au}R^{u}_{bc}-g_{bc}\frac{\partial f}{\partial q^{a}}-g_{ac}\frac{\partial f}{\partial q^{b}}+2g_{ab}\frac{\partial f}{\partial q^{c}}.$
@@
We conclude that, despite the fact that the system is not $\varphi$-simple, we do have a projective geodesic extension that is a $\mathcal{D}$-comformal modification of $g$. Indeed, we have that $(F=f\circ\pi,\overline{g}_{ai}=-G_{ai}=-1)$, for the $f$ above, is such a projective geodesic extension. The nonholonomic trajectories are pregeodesics of the metric defined by the Lagrangian
@@
is positive definite. The characteristic polynomial of this matrix is given by $(1-\lambda)^{3}(\alpha+2\lambda-3)=0$. This means that there exists a (strictly positive) eigenvalue $\lambda=1$, with multiplicity 3. The other eigenvalue is $\frac{3-\alpha}{2}$. If we want it to be strictly larger than zero, we may take e.g. $\alpha=\overline{g}_{uu}=-42$.

### v2

@@ 6.4 A mathematical example
Based on the classification in Proposition 1 we know that $(A^{\prime})^{G}$ is a weaker condition than the $\varphi$-simplicity property, but a stronger condition than having only an invariant volume form. In [20] it is shown that the ball rolling over a sphere is an example of a Chaplygin system that possesses an invariant measure, but that is not Hamiltonizable, and thus also not $\varphi$-simple. To further verify that the item $(ii)$ of the classification is not void, we give a (mathematical) example of a nonholonomic system that is not $\varphi$-simple, but that does have a solution $f$ of condition $(A^{\prime})^{G}$. Such a system will allow for a projective geodesic extension (with $\overline{g}_{ai}=-G_{ai}$) and an invariant volume form, without being $\varphi$-simple.
@@
$=$ $\frac{1}{2}({\tilde{v}_{x}}^{2}+{\tilde{v}_{y}}^{2}+{\tilde{v}_{z}}^{2}+2{\tilde{v}_{x}}{\tilde{v}_{u}}+2{\tilde{v}_{y}}{\tilde{v}_{u}}+2{\tilde{v}_{z}}{\tilde{v}_{u}}+4{\tilde{v}_{u}}^{2}),$
@@
$G_{bu}R_{ac}^{u}-G_{au}R^{u}_{bc}-g_{bc}\frac{\partial f}{\partial q^{a}}-g_{ac}\frac{\partial f}{\partial q^{b}}+2g_{ab}\frac{\partial f}{\partial q^{c}}=0.$
@@
We conclude that, despite the fact that the system is not $\varphi$-simple, we do have a projective geodesic extension that is a $\mathcal{D}$-conformal modification of $g$. Indeed, we have that $(F=f\circ\pi,\overline{g}_{ai}=-G_{ai}=-1)$, for the $f$ above, is such a projective geodesic extension. The nonholonomic trajectories are pregeodesics of the metric defined by the Lagrangian
@@
is positive definite. If we set $a:=\frac{1}{2}e^{2h(x+y+z)}$, then the characteristic polynomial of this matrix is given by $(a-\lambda)^{2}(a^{2}\alpha-3a^{2}+(-a\alpha-a)\lambda+\lambda^{2})=0$. This means that there exists a (strictly positive) eigenvalue $\lambda=a$, with multiplicity 2. The two other eigenvalues have a sum of $a(\alpha+1)$ and product $a^{2}(\alpha-3)$. We can ensure that they are both strictly larger than zero by taking e.g. $\alpha=\overline{g}_{uu}=42$.
@@ 7 Hamiltonian reparametrization by geodesic equations
Acknowledgements. We thank the Research Fund of the University of Antwerp (BOF) for its support through the DOCPRO project 49747.